Brandon is 6 times as old as Cora. In 4 years, Brandon will be only twice as old as Cora will be then. Find Brandon's age now.
Answer by richwmiller(17219)

b=6c
b+4=2(c+4)
6c+4=2(c+4)
6c+4=2c+8
4c=4
c=1
b=6